# Revisiting Convergence of AdaGrad with Relaxed Assumptions

### v1

@@ 5.1 Preliminary
We let the function value gap $\Delta^{(y)}_{s}=f(\bm{y}_{s})-f^{*}$ and introduce $\{\mathcal{G}_{s}\}_{s\geq 1}$ and the value $\mathcal{G}$,
$\begin{split}&\mathcal{G}_{s}=\sqrt{X\Delta^{(x)}_{s}+2C},\quad\mathcal{G}=\sqrt{X\Delta+2C},\quad X=2A+4LB+4L,\end{split}$ (9)
@@ 5.2 Rough estimations
$\sum_{t=1}^{T}$ $\Delta^{(x)}_{t}\leq\Delta^{(x)}_{1}T+\left(\frac{\eta\|\bar{\bm{g}}_{1}\|\sqrt{d}}{1-\beta}+\frac{L\eta^{2}d}{2(1-\beta)^{2}}\right)T^{2}+\frac{L\eta^{2}dT^{3}}{(1-\beta)^{2}}.$
@@ 5.3 Start Point and decomposition
We now proceed the proof for the main result. We fix the horizon $T$. Following [Ward et al., 2020], we start from the descent lemma of smoothness over $\bm{y}_{s}$ with both sides subtracting $f^{*}$,
@@
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}+\frac{\eta}{1-\beta}\underbrace{\left(-\sum_{s=1}^{t}\left\langle\nabla f(\bm{y}_{s}),\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\rangle\right)}_{\textbf{A}}+\frac{L\eta^{2}}{2(1-\beta)^{2}}\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2},$ (10)
@@ 5.4 Estimating A
${\bf A.1}=-\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}\underbrace{-\sum_{s=1}^{t}\left\langle\bar{\bm{g}}_{s},\frac{\bm{\xi}_{s}}{\bm{a}_{s}}\right\rangle}_{{\bf A.1.1}}+\underbrace{\sum_{s=1}^{t}\left\langle\bar{\bm{g}}_{s},\left(\frac{1}{\bm{a}_{s}}-\frac{1}{\bm{b}_{s}}\right)\bm{g}_{s}\right\rangle}_{\textbf{A.1.2}}.$ (13)
@@ 5.5 Bounding the function value gap
Based on the above estimations, we could use an induction argument to deduce an upper bound for function value gaps.
@@
$\Delta^{(x)}_{t+1}\leq{\Delta}-\frac{\eta}{1-\beta}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2},\quad\forall t\in[T].$ (18)
@@
$\frac{\Delta^{(x)}_{t+1}}{2}\leq\frac{\Delta}{4}+\Delta^{(x)}_{1}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}+\frac{L\|\bm{m}_{t}\|^{2}}{2(1-\beta)^{2}}+\frac{3\sqrt{2C}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)$
$+\left(\frac{\sqrt{2C}\eta}{1-\beta}+\tilde{L}\right)\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}+\sum_{s=1}^{t}\frac{L\|\bm{m}_{s-1}\|^{2}}{2(1-\beta)}+\frac{18X\eta^{2}}{(1-\beta)^{2}}\log^{2}\left(\frac{T}{\delta}\right)+\frac{2X\eta^{2}}{(1-\beta)^{2}}\left(\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}\right)^{2}.$ (22)
@@
$\sum_{s=1}^{t}\|\bm{m}_{s}\|^{2}\leq\frac{\eta^{2}d}{(1-\beta)^{2}}\log\mathcal{F}_{T}$
@@
Compared with Lemma 5.1, Lemma 5.8 improves the dependency of $1-\beta$ for estimating $\|\bm{m}_{t}\|$, which leads to the $\mathcal{O}((1-\beta)^{-1})$ order for the final convergence as in Remark 4.1. Thus, applying Lemma 5.8 over (22), and then combining with $\Delta$ in (20),
@@ 5.6 Proof of the main result
$\quad\sum_{s=1}^{T}\|\bar{\bm{g}}_{s}\|^{2}-\Delta_{1}\epsilon\leq\Delta_{1}\left(\sqrt{2B_{1}\sum_{s=1}^{T}\|\bar{\bm{g}}_{s}\|^{2}}+\sqrt{2(A\Delta+C)T}+\mathcal{G}\right)$
@@ 6.2 Convergence result
Let $T\geq 1$ and $\delta\in(0,1)$. Suppose that $\{\bm{x}_{s}\}_{s\in[T]}$ is a sequence generated by Algorithm 1, $f$ is $(L_{0},L_{1})$-smooth satisfying (25), Assumptions (A1), (A2), (A3) hold, and the parameters satisfy that $\beta\in[0,1)$,
@@
$\begin{split}&\mathcal{H}=\sqrt{(2A+4LB+4L)\Lambda_{x}+2C},\quad\mathcal{L}=2L_{0}+2L_{1}\left(4L_{1}\Lambda_{x}+\sqrt{4L_{0}\Lambda_{x}}\right),\end{split}$ (28)
@@
$\Lambda_{y}\sim\mathcal{O}\left(\Delta^{(x)}_{1}+\frac{C_{0}^{2}d+C_{0}d}{(1-\beta)^{3}}\log\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right),\quad\tilde{\Lambda}_{y}=\frac{2\Lambda_{y}(1-\beta)}{\eta},\quad\Lambda_{x}\sim\mathcal{O}\left(\Lambda_{y}^{2}\right).$
@@
$\frac{1}{T}\sum_{s=1}^{T}\|\nabla f(\bm{x}_{s})\|^{2}\leq\mathcal{O}\left(\tilde{\Lambda}_{y}\left(\frac{B_{1}\tilde{\Lambda}_{y}+\sqrt{B_{1}L\Lambda_{x}}+\epsilon}{T}+\sqrt{\frac{A\Lambda_{x}+C}{T}}\right)\right).$
It’s easy to verify that $\Lambda_{y},\tilde{\Lambda}_{y},\Lambda_{x},\mathcal{H},\mathcal{L}$ are all restricted by polynomials of $\log T$. Then from (27), when $T\gg d$, a typical setting is $\eta\sim 1/\mathcal{O}(\log^{2}(T/\delta))$. Moreover, the convergence rate is still adaptive to the noise parameters $A,C$ and requires problem-parameters to tune step-sizes potentially due to the relaxation of smoothness. The subsequent result for AdaGrad under (25) could be directly deduced from Theorem 2 and will be presented in Appendix.
@@ Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under grant number 2021YFA1003500. The corresponding author is Junhong Lin.
@@ Appendix A Complementary lemmas
Then for any $\lambda>0$, and for any $\delta\in(0,1)$, it holds that
@@ Appendix B Omitted proofs under smooth case
$\mathbb{P}\left(\max_{s\in[T]}\gamma_{s}\geq E\right)=\mathbb{P}\left(\exp\left(\max_{s\in[T]}\gamma_{s}\right)\geq\mathrm{e}^{E}\right)$
$\leq$ $\mathrm{e}^{-E}\mathbb{E}\left[\exp\left(\max_{s\in[T]}\gamma_{s}\right)\right]\leq\mathrm{e}^{-E}\mathbb{E}\left[\sum_{s=1}^{T}\exp\left(\gamma_{s}\right)\right]\leq\mathrm{e}^{-E}T\mathrm{e},$
@@
Compared with those under Assumption (A3), an additional $\log T$ factor emerges in (29). Hence, $\mathcal{G}_{s}$ and $\mathcal{G}$ defined in (9) should be revised as
@@
Noting that $|g_{s,i}|\leq\sqrt{v_{s,i}},\forall i\in[d]$. We therefore verify that $\|\bm{g}_{s}/\sqrt{\bm{v}_{s}}\|\leq\sqrt{d}\max_{i\in[d]}|g_{s,i}/v_{s,i}|\leq\sqrt{d}$. Moreover,
@@
Combining the above, we obtain that
@@
$\Delta^{(x)}_{t+1}$ $\leq\Delta^{(x)}_{1}+\frac{\eta\sqrt{d}}{1-\beta}\sum_{s=1}^{t}\left(\|\bar{\bm{g}}_{1}\|+\frac{L\eta\sqrt{d}s}{1-\beta}\right)+\frac{L\eta^{2}dt}{2(1-\beta)^{2}}$
$\leq\Delta^{(x)}_{1}+\left(\frac{\eta\|\bar{\bm{g}}_{1}\|\sqrt{d}}{1-\beta}+\frac{L\eta^{2}d}{2(1-\beta)^{2}}\right)t+\frac{L\eta^{2}dt^{2}}{(1-\beta)^{2}}.$
Finally, we sum up over $t\in[T]$ and obtain that
$\sum_{t=1}^{T}\Delta^{(x)}_{t}$ $\leq\sum_{t=1}^{T}\Delta^{(x)}_{1}+\left(\frac{\eta\|\bar{\bm{g}}_{1}\|\sqrt{d}}{1-\beta}+\frac{L\eta^{2}d}{2(1-\beta)^{2}}\right)\sum_{t=1}^{T}(t-1)+\frac{L\eta^{2}d}{(1-\beta)^{2}}\sum_{t=1}^{T}(t-1)^{2}$
@@
Thus, $\|\bm{g}_{s}\|\leq\mathcal{G}_{s},\forall s\geq 1$. Let $\bm{a}_{s}=\sqrt{\tilde{\bm{v}}_{s}}+\bm{\epsilon}$ where $\tilde{\bm{v}}_{s}=(\tilde{v}_{s,i})_{i}$. Then, for any $i\in[d]$,
$\left|\frac{1}{a_{s,i}}-\frac{1}{b_{s,i}}\right|$ $=\frac{|\sqrt{\tilde{v}_{s,i}}-\sqrt{v_{s,i}}|}{a_{s,i}b_{s,i}}=\frac{|\tilde{v}_{s,i}-v_{s,i}|}{a_{s,i}b_{s,i}(\sqrt{\tilde{v}_{s,i}}+\sqrt{v_{s,i}})}$
$=\frac{1}{a_{s,i}b_{s,i}}\frac{|\mathcal{G}_{s}^{2}-g_{s,i}^{2}|}{\sqrt{v_{s,i}}+\sqrt{\tilde{v}_{s,i}}}\leq\frac{1}{a_{s,i}b_{s,i}}\frac{\mathcal{G}_{s}^{2}}{\sqrt{v_{s,i}}+\sqrt{\tilde{v}_{s,i}}}\leq\frac{\mathcal{G}_{s}}{a_{s,i}b_{s,i}}.$ (35)
@@
${\bf A.1.2}$ $\leq\sum_{s=1}^{t}\sum_{i=1}^{d}\left|\frac{1}{a_{s,i}}-\frac{1}{b_{s,i}}\right||\bar{g}_{s,i}||g_{s,i}|\leq\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\mathcal{G}_{s}}{a_{s,i}b_{s,i}}|\bar{g}_{s,i}||g_{s,i}|$
$\leq\frac{1}{4}\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\bar{g}_{s,i}^{2}}{a_{s,i}}+\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\mathcal{G}_{s}^{2}}{a_{s,i}}\frac{g_{s,i}^{2}}{b_{s,i}^{2}}\leq\frac{1}{4}\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\bar{g}_{s,i}^{2}}{a_{s,i}}+\sum_{s=1}^{t}\sum_{i=1}^{d}{\mathcal{G}_{s}}\frac{g_{s,i}^{2}}{b_{s,i}^{2}},$ (36)
@@
Finally, we could follow the similar analysis for (38) and (39) to deduce that the terms inside the logarithm operator could be further bounded by $\mathcal{F}_{T}$ and thereby verify the target results. ∎

### v2

@@ 5.1 Preliminary
We let the function value gap $\Delta^{(y)}_{s}=f(\bm{y}_{s})-f^{*}$. In addition, we introduce $\{\mathcal{G}_{s}\}_{s\geq 1}$ and the value $\mathcal{G}$,
$\begin{split}&\mathcal{G}_{s}=\sqrt{X\Delta^{(x)}_{s}+2C},\\ &\mathcal{G}=\sqrt{X\Delta+2C},\quad X=2A+4LB+4L,\end{split}$ (9)
@@ 5.2 Rough estimations
$\sum_{t=1}^{T}$ $\Delta^{(x)}_{t}\leq\Delta^{(x)}_{1}T$
$+\left(\frac{\eta\|\bar{\bm{g}}_{1}\|\sqrt{d}}{1-\beta}+\frac{L\eta^{2}d}{2(1-\beta)^{2}}\right)T^{2}+\frac{L\eta^{2}dT^{3}}{(1-\beta)^{2}}.$
@@ 5.3 Start Point and decomposition
We now proceed the proof for the main result. We fix the horizon $T$. Following [Ward et al., 2020], we start from the descent lemma of smoothness over $\bm{y}_{s}$ with both sides subtracting with $f^{*}$,
@@
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}+\frac{\eta}{1-\beta}\underbrace{\left(-\sum_{s=1}^{t}\left\langle\nabla f(\bm{y}_{s}),\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\rangle\right)}_{\textbf{A}}$
$\quad+\frac{L\eta^{2}}{2(1-\beta)^{2}}\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2},$ (10)
@@ 5.4 Estimating A
${\bf A.1}=-\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}$
$\quad\quad\underbrace{-\sum_{s=1}^{t}\left\langle\bar{\bm{g}}_{s},\frac{\bm{\xi}_{s}}{\bm{a}_{s}}\right\rangle}_{{\bf A.1.1}}+\underbrace{\sum_{s=1}^{t}\left\langle\bar{\bm{g}}_{s},\left(\frac{1}{\bm{a}_{s}}-\frac{1}{\bm{b}_{s}}\right)\bm{g}_{s}\right\rangle}_{\textbf{A.1.2}}.$ (13)
@@ 5.5 Bounding the function value gap
Based on the above estimations, we could use an induction argument to deduce an upper bound for function value gaps. The induction technique is motivated by [Attia and Koren, 2023] where AdaGrad-Norm with affine variance noise was studied. As we study a more relaxed assumption on AdaGrad, it’s required to provide some new estimations.
@@
$\Delta^{(x)}_{t+1}\leq{\Delta}-\frac{\eta}{1-\beta}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2},\quad\forall t\in[T],$ (18)
@@
$\frac{\Delta^{(x)}_{t+1}}{2}\leq\frac{\Delta}{4}+\Delta^{(x)}_{1}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}$
$+\frac{L\|\bm{m}_{t}\|^{2}}{2(1-\beta)^{2}}+\frac{3\sqrt{2C}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)$
$+\left(\frac{\sqrt{2C}\eta}{1-\beta}+\tilde{L}\right)\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}+\sum_{s=1}^{t}\frac{L\|\bm{m}_{s-1}\|^{2}}{2(1-\beta)}$
$+\frac{18X\eta^{2}}{(1-\beta)^{2}}\log^{2}\left(\frac{T}{\delta}\right)+\frac{2X\eta^{2}}{(1-\beta)^{2}}\left(\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}\right)^{2}.$ (22)
@@
$\sum_{s=1}^{t}\|\bm{m}_{s}\|^{2}\leq\frac{\eta^{2}d}{(1-\beta)^{2}}\log\mathcal{F}_{T},$
@@
Compared with Lemma 5.1, Lemma 5.8 improves the dependency to $1-\beta$ for estimating $\|\bm{m}_{t}\|^{2}$, which leads to the $\mathcal{O}((1-\beta)^{-1})$ order for the final convergence as in Remark 4.1. Thus, applying Lemma 5.8 over (22), and then combining with $\Delta$ in (20),
@@ 5.6 Proof of the main result
$\quad\sum_{s=1}^{T}\|\bar{\bm{g}}_{s}\|^{2}-\Delta_{1}\epsilon$
$\leq\Delta_{1}\left(\sqrt{2B_{1}\sum_{s=1}^{T}\|\bar{\bm{g}}_{s}\|^{2}}+\sqrt{2(A\Delta+C)T}+\mathcal{G}\right)$
@@ 6.2 Convergence result
Let $T\geq 1$ and $\delta\in(0,1)$. Suppose that $\{\bm{x}_{s}\}_{s\in[T]}$ is a sequence generated by Algorithm 1, $f$ is $(L_{0},L_{1})$-smooth satisfying (25), Assumptions (A1), (A2), (A3) hold, and the hyper-parameters satisfy that $\beta\in[0,1)$,
@@
$\begin{split}&\mathcal{H}=\sqrt{2A\Lambda_{x}+2(B+1)\left(4L_{1}\Lambda_{x}+\sqrt{4L_{0}\Lambda_{x}}\right)^{2}+2C},\\ &\mathcal{L}=2L_{0}+2L_{1}\left(4L_{1}\Lambda_{x}+\sqrt{4L_{0}\Lambda_{x}}\right),\end{split}$ (28)
@@
$\Lambda_{y}\sim\mathcal{O}\left(\Delta^{(x)}_{1}+\frac{C_{0}^{2}d+C_{0}d}{(1-\beta)^{3}}\log\left(\frac{T}{\delta}+\frac{T}{\epsilon^{2}}\right)\right),$
$\tilde{\Lambda}_{y}=\frac{2\Lambda_{y}(1-\beta)}{\eta},\quad\Lambda_{x}\sim\mathcal{O}\left(\Lambda_{y}^{2}\right).$
@@
$\frac{1}{T}\sum_{s=1}^{T}\|\nabla f(\bm{x}_{s})\|^{2}$
$\leq$ $\mathcal{O}\left(\tilde{\Lambda}_{y}\left(\frac{B_{1}\tilde{\Lambda}_{y}+\sqrt{B_{1}L\Lambda_{x}}+\epsilon}{T}+\sqrt{\frac{A\Lambda_{x}+C}{T}}\right)\right).$
It’s easy to verify that $\Lambda_{y}\sim\mathcal{O}(\log(T/\delta))$ and thereby $\Lambda_{x},\mathcal{H},\mathcal{L}\sim\mathcal{O}(\log^{2}(T/\delta))$. Then, from (27), when $T\gg d$, a typical setting is $\eta\sim\mathcal{O}(\log^{2}(T/\delta))$. Moreover, the convergence rate is still adaptive to the noise parameters $A,C$ and requires problem-parameters to tune step-sizes potentially due to the relaxation of smoothness. The subsequent result for AdaGrad under (25) could be directly deduced from Theorem 2 and will be presented in Appendix.
7 Conclusion
In this paper, we provide high probability convergence bounds for AdaGrad and its momentum variant under the non-convex smooth optimization. In particular, we consider a mild noise model incorporating affine variance noise and the expected smoothness. We rely on a new proxy step-size and some delicate estimations to derive the bound. Our findings reveal that without fined-tuning step-sizes by the smooth parameter and noise level, AdaGrad can find a stationary point with a rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$, particularly accelerating to $\tilde{\mathcal{O}}(1/T)$ when specific noise parameters are sufficiently small. Furthermore, we extend our framework to the generalized smooth case that allows for unbounded smooth parameters, showing the same convergence rate, albeit that fined-tuning step-sizes are required in the latter.
Limitation
Although AdaGrad plays an important role in the adaptive method field, several other adaptive methods including RMSProp, Adam and AdamW, may be preferred in some real applications. Therefore, it is also pertinent to study these algorithms under relaxed assumptions. In addition, it is still unknown whether similar convergence result could be also achieved under an expected version of Assumption (A3). Finally, as we study a new assumption over AdaGrad, it would be more beneficial to provide more experimental results to support the theoretical results.
@@ Acknowledgments
The authors would like to thank the reviewers and area chairs for their constructive comments. This work was supported in part by the National Key Research and Development Program of China under grant number 2021YFA1003500, and NSFC under grant numbers 11971427. We also thank Chenhao Yu very much on pointing out one error in the generalized-smooth analysis, and Mouxiang Chen for his great help with the experimental results. The corresponding author is Junhong Lin.
@@ Appendix A Complementary lemmas
Then, for any $\lambda>0$, and for any $\delta\in(0,1)$, it holds that
@@ Appendix B Omitted proofs under smooth case
$\mathbb{P}\left(\max_{s\in[T]}\gamma_{s}\geq E\right)$ $=\mathbb{P}\left(\exp\left(\max_{s\in[T]}\gamma_{s}\right)\geq\mathrm{e}^{E}\right)\leq\mathrm{e}^{-E}\mathbb{E}\left[\exp\left(\max_{s\in[T]}\gamma_{s}\right)\right]\leq\mathrm{e}^{-E}\mathbb{E}\left[\sum_{s=1}^{T}\exp\left(\gamma_{s}\right)\right]\leq\mathrm{e}^{-E}T\mathrm{e},$
@@
Compared (29) with Assumption (A3), an additional $\log T$ factor emerges. Hence, $\mathcal{G}_{s}$ and $\mathcal{G}$ defined in (9) should be revised as
@@
Note that $|g_{s,i}|\leq\sqrt{v_{s,i}},\forall i\in[d]$. We therefore verify that $\|\bm{g}_{s}/\sqrt{\bm{v}_{s}}\|\leq\sqrt{d}\max_{i\in[d]}|g_{s,i}/v_{s,i}|\leq\sqrt{d}$. Moreover,
@@
Combining with the above, we obtain that
@@
$\Delta^{(x)}_{t+1}$ $\leq\Delta^{(x)}_{1}+\frac{\eta\sqrt{d}}{1-\beta}\sum_{s=1}^{t}\left(\|\bar{\bm{g}}_{1}\|+\frac{L\eta\sqrt{d}s}{1-\beta}\right)+\frac{L\eta^{2}dt}{2(1-\beta)^{2}}.$ (33)
We define $\sum_{a}^{b}=0$ when $a<b$. Then, we sum up both sides of (33) over $t\in[0,1,\cdots,T-1]$ to obtain that
$\sum_{t=1}^{T}\Delta^{(x)}_{t}$ $\leq\sum_{t=0}^{T-1}\Delta^{(x)}_{1}+\frac{\eta\sqrt{d}}{1-\beta}\sum_{t=0}^{T-1}\sum_{s=1}^{t}\left(\|\bar{\bm{g}}_{1}\|+\frac{L\eta\sqrt{d}s}{1-\beta}\right)+\sum_{t=0}^{T-1}\frac{L\eta^{2}dt}{2(1-\beta)^{2}}$
$\leq\Delta^{(x)}_{1}T+\frac{\eta\sqrt{d}}{1-\beta}\sum_{t=1}^{T}\sum_{s=1}^{t}\left(\|\bar{\bm{g}}_{1}\|+\frac{L\eta\sqrt{d}s}{1-\beta}\right)+\sum_{t=1}^{T}\frac{L\eta^{2}dt}{2(1-\beta)^{2}}$
@@
Thus, $\|\bm{g}_{s}\|\leq\mathcal{G}_{s},\forall s\geq 1$. Let $\bm{a}_{s}=\sqrt{\tilde{\bm{v}}_{s}}+\bm{\epsilon}$ where $\tilde{\bm{v}}_{s}=(\tilde{v}_{s,i})_{i}$. Note that $\tilde{v}_{s,i}\geq\mathcal{G}_{s}$. Then, for any $i\in[d]$,
$\left|\frac{1}{a_{s,i}}-\frac{1}{b_{s,i}}\right|=\frac{|\sqrt{\tilde{v}_{s,i}}-\sqrt{v_{s,i}}|}{a_{s,i}b_{s,i}}=\frac{|\tilde{v}_{s,i}-v_{s,i}|}{a_{s,i}b_{s,i}(\sqrt{\tilde{v}_{s,i}}+\sqrt{v_{s,i}})}=\frac{1}{a_{s,i}b_{s,i}}\frac{|\mathcal{G}_{s}^{2}-g_{s,i}^{2}|}{\sqrt{v_{s,i}}+\sqrt{\tilde{v}_{s,i}}}$
$\leq$ $\frac{1}{a_{s,i}b_{s,i}}\frac{\mathcal{G}_{s}^{2}}{\sqrt{v_{s,i}}+\sqrt{\tilde{v}_{s,i}}}\leq\frac{\mathcal{G}_{s}}{a_{s,i}b_{s,i}}.$ (36)
@@
${\bf A.1.2}$ $\leq\sum_{s=1}^{t}\sum_{i=1}^{d}\left|\frac{1}{a_{s,i}}-\frac{1}{b_{s,i}}\right||\bar{g}_{s,i}||g_{s,i}|\leq\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\mathcal{G}_{s}}{a_{s,i}b_{s,i}}|\bar{g}_{s,i}||g_{s,i}|\leq\frac{1}{4}\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\bar{g}_{s,i}^{2}}{a_{s,i}}+\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\mathcal{G}_{s}^{2}}{a_{s,i}}\frac{g_{s,i}^{2}}{b_{s,i}^{2}}$
$\leq\frac{1}{4}\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\bar{g}_{s,i}^{2}}{a_{s,i}}+\sum_{s=1}^{t}\sum_{i=1}^{d}{\mathcal{G}_{s}}\frac{g_{s,i}^{2}}{b_{s,i}^{2}},$ (37)
@@
Finally, we could follow the similar analysis for (39) and (40) to deduce that the terms inside the logarithm operator in (44) could be further bounded by $\mathcal{F}_{T}$ and thereby verify the target results. ∎